The information necessary for preparing the 2021 year-end adjusting entries for Gamecock Advertising Agency appears below. Gamecock’s fiscal year-end is December 31.

  1. On July 1, 2021, Gamecock receives $4,500 from a customer for advertising services to be given evenly over the next 10 months. Gamecock credits Deferred Revenue.
  2. At the beginning of the year, Gamecock’s depreciable equipment has a cost of $27,500, a five-year life, and no salvage value. The equipment is depreciated evenly (straight-line depreciation method) over the five years.
  3. On May 1, 2021, the company pays $3,000 for a two-year fire and liability insurance policy and debits Prepaid Insurance.
  4. On September 1, 2021, the company borrows $25,000 from a local bank and signs a note. Principal and interest at 9% will be paid on August 31, 2022.
  5. At year-end there is a $1,950 debit balance in the Supplies (asset) account. Only $850 of supplies remains on hand.

      
Required:

Record the necessary adjusting entries on December 31, 2021. No prior adjustments have been made during 2021. (Do not round intermediate calculations. If no entry is required for a particular transaction/event, select "No Journal Entry Required" in the first account field.)
Answer #1
Debit Credit
1 Deferred Revenue 2700 =4500*6/10
     Service revenue 2700
2 Depreciation expense 5500 =27500/5
       Accumulated depreciation 5500
3 Insurance expense 1000 =3000*8/24
     Prepaid insurance 1000
4 Interest expense 750 =25000*9%*4/12
      Interest payable 750
5 Supplies expense 1100 =1950-850
      Supplies 1100